a plane is flying from NYC to SF. It normally flies 300 mph. there is 30 mph wind heading west to east. How long does it take to fly 200 miles to SF and how long does it take to fly 200 miles back to NYC
Answer by checkley71(8403)

D=RT DISTANCE=RATE*TIME
200=(300-30)T
200=270T
270T=200
T=200/270
T=.74NHOURS TO GO FROM NYC TO SF.
200=(300+30)T
200=330T
330T=200
T=200/330
T=.6 HOURS FOR THE RETURN TRIP.
---------------------------------------------------------------
IF THE DISTANCE IS SUPPOSED TO BE 2000 MILES FROM NYC TO SAN FRANCISCO.
JUST MULTIPLY THE ANSWERS OF .74 BY 10=7.4 HOURS & THE RETURN TRIP TAKES .6*10=6 HOURS.